OpenStudy (anonymous):

I need help /.\ A basketball is thrown upwards. The height f(t), in feet, of the basketball at time t, in seconds, is given by the following function: f(t) = −16t^2 + 16t + 32 Which of the following is a reasonable domain of the graph of the function when the basketball falls from its maximum height to the ground? 0.5 < t < 1 0.5 < t < 2 1 < t < 2 1 < t < 1.5

OpenStudy (anonymous):

I'm thinking 0.5 < t < 2

OpenStudy (anonymous):

I'm just confused, fully with the entire thing. If you don't mind, could you help me with another question?

OpenStudy (anonymous):

highest point should be the vertex and the rightest zero x value of vertex = -b/2a given ax^2 + bx + c so plug in and you will get your first point

OpenStudy (anonymous):

the other point would be solve -16x^2 + 16x + 32 = 0 and take the positive answer, then you will get your other point

OpenStudy (anonymous):

Ohhh okay. I have one more question that I don't get, can you help me?

OpenStudy (anonymous):

Yes I will help with what I can :)

OpenStudy (anonymous):

Thank you ^.^ A punter kicks a football upward with an initial speed of 48 feet per second. After how many seconds does the ball hit the ground? Use the formula h = rt − 16t2 where h represents height in feet and r represents the initial speed.

OpenStudy (anonymous):

h = 0 when it hits the ground, therefore: -16t^2 + 48t = 0 _________________________ Factor out -16t -16t(t + 3) = 0 : Two solutions t = 0, when it begins it's upward journey and t = 3 sec, when it returns to earth

OpenStudy (anonymous):

So the answer would end up being 3?

OpenStudy (anonymous):

Yes :)

OpenStudy (anonymous):

Okay, thank you so much! That last one was easier than I thought, I thought there were more steps cx

OpenStudy (anonymous):

No problem :P